OpenStudy (anonymous):

how many positive integers less than 220 have a remainder of 3 when divided by 9?

OpenStudy (anonymous):

first lets start by listing some of the integers that have a remainder of 3 when divided by 9. 3, 12, 21, 30, 39,...., 219 So we just need to figure out how many numbers are in that list.

OpenStudy (anonymous):

All of these numbers are in the form: 9k+3, for k =0, 1, 2, 3, 4.... we can see this because: 9(0)+3 = 3 9(1)+3 = 12 so on and so forth.

OpenStudy (anonymous):

Just find the last of the possible ones which is 219, and as stated previously in the form 9k+3, therefore 219=9k+3 , k=24 , hence your answer is 24 integers.

OpenStudy (anonymous):

so to figure out how many numbers there are, we solve: 219 = 9k+3 216 = 9k 24 = k so there are 25 numbers on the list.

OpenStudy (anonymous):

The equation will be 9A+3 Put the value of A = 0,1,2,3....24 Get the series .. Apply A.P get n=?

OpenStudy (anonymous):

25

OpenStudy (anonymous):

its 25 numbers, because we started at k = 0

OpenStudy (anonymous):

My apologies, 3 is also included , so it is 25 integers